Given the demand for a product as Qd=26-8P and the supply as Qs=-16+6P. You are told equilibrium is obtained at the point where Qd=As.The equilibrium price for the product is


1
Expert's answer
2022-04-24T17:24:41-0400

"Q_d=26-8P,"

"Q_s=-16+6P,"

"Q_d=Q_s,"

"26-8P=-16+6P,"

"42=14P,"

"P=3,"

"Q_d=Q_s=2."